OpenStudy (anonymous):

How do I write [4(cos(7π/9)+isin(7π/9))]^3 in standard form for the complex number a+bi.

OpenStudy (anonymous):

\(\large [4(cos(7π/9)+isin(7π/9))]^3 \)

OpenStudy (anonymous):

hint 1:- (cos theta +i sin theta ) ^n=(cos n theta +i sin n theta ) hint 2:- e^itheta =(cos theta +i sin theta )

OpenStudy (anonymous):

First use Demoivre's theorem to get rid of the exponent: \[\large \left(re^{i\theta}\right)^n=r^ne^{in\theta}\] So, \[\bigg[4\left(\cos\frac{7\pi}{9}+i\sin\frac{7\pi}{9}\right)\bigg]^3=64\left(\cos\frac{7\pi}{3}+i\sin\frac{7\pi}{3}\right)\] Find the trig ratios and simplify.

OpenStudy (anonymous):

OK thanks